perimeter is 220 feet, the length of the field is 2 feet more than the width, find the length
Answer by ikleyn(52946)

.

Perimeter = 2L + 2*(L-2) = 220


4L - 4 = 220

4L = 224

L = 224/4 = 56 feet.    ANSWER


CHECK.  2*56 + 2*54 = 2*110 = 220.   ! Correct !

Solved.